40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Jan 30 21:29:36 2018, awade, Craig, DailyProgress, PMC, Pre mode cleaner installed in North path 
    Reply  Thu Feb 1 15:25:58 2018, awade, Craig, DailyProgress, PMC, Update working schematic of PMC servo card (older card) D980352-A_AWade20180201Mods.pdf
       Reply  Sat Feb 3 18:08:26 2018, awade, DailyProgress, PMC, A few more modifications to the PMC north board D980352-A_AWade20180202Mods.pdf
          Reply  Fri Jul 6 23:43:54 2018, anchal, DailyProgress, PMC, New modifications to the PMC north board LowPrefFilter.filPMCLowPrefFilterOpt.ipynbU2Buffer.filU2Stability.ipynbD980352-A_AWade20180201Mods_--AGupta20180707Mods.pdf
             Reply  Mon Jul 9 16:15:06 2018, rana, DailyProgress, PMC, New modifications to the PMC north board 
                Reply  Thu Jul 12 15:18:26 2018, anchal, DailyProgress, PMC, New modifications to the PMC north board D980352-AGupta20180712Mods.pdf
                   Reply  Sat Jul 21 14:53:16 2018, awade, DailyProgress, PMC, New modifications to the PMC north board 
                      Reply  Wed Aug 29 12:01:54 2018, anchal, Summary, PMC, Summarization of efforts to fix PMC Servo problems D980352-AGupta20180712Mods.pdfPMC_Servo_LISO_Analysis_Vs_Measured.pdf
                         Reply  Thu Aug 30 12:01:37 2018, rana, Summary, PMC, Summarization of efforts to fix PMC Servo problems 
                         Reply  Mon Sep 24 12:53:26 2018, awade and anchal, Summary, PMC, Summarization of efforts to fix PMC Servo problems: external mixer and LP filter 
Message ID: 2070     Entry time: Sat Feb 3 18:08:26 2018     In reply to: 2066     Reply to this: 2209
Author: awade 
Type: DailyProgress 
Category: PMC 
Subject: A few more modifications to the PMC north board 

A few more changes to the PMC north board. Hopefully this is the last.

Changing series resistance at the output

There was some limit to the gain before the PMC loop became unstable.  I realized that I hadn't implemented the 10 Hz pole at the output, that could be the reason.  The resistor in series with the output was only 1 kΩ.  The Noliac NAC2124 PZT on the PMC cavity has a specified capacitance of 510 nF and about 5 meters of BNC cabling should have a capacitance of 0.5 nF (for 101pF/m).  This should give a pole of about 312 Hz, we want to bring this down.  I replaced the 1 kΩ series resistance at the HV output with a 30.9 kΩ to bring to pole down to 10 Hz.  

Schematic is updated below.

Removing short from PD RF input

For whatever reason the AC coupling resistor had been deliberately shorted on the PD RF input to the board (C3).  I removed the short.  It is AC coupled again.

 

Altering the Acromag excitation voltage and divider resistors

I realized running the Acromags off 6 V is a little awkward.  Most power supplies come in other standard values like 5 V, 7.5 V, 9V etc.  I've raised the excitation of the binary channels to 9 V and made some changes to the divider resistors. Resistors on SW1 and SW2 binary engages were changed to equal values of 1790 Ω for R1 and R2 (see option D, attachment, PSL:2058). On the blanking pin used to activate the AD602, equal values for R1 and R2 of 154 Ω were used.  These resistor choices give a high voltage of 4.5 V and a low of 0.66 V.  There is also a margin of error if the power supply used goes a little over voltage.

I've ordered standard barrel 2.1x5.5 mm panel Jacks so that the unit can be powered from a standard 9V plug pack.

Attachment 1: D980352-A_AWade20180202Mods.pdf  309 kB  | Hide | Hide all
D980352-A_AWade20180202Mods.pdf
ELOG V3.1.3-